A 1.5 m of pendulum is swinging on the planet Jupiter with a period of 3.7 s, calculate the value of (g) on the planet Jupiter.


1
Expert's answer
2021-01-26T19:27:05-0500

By the definition of the period of the pendulum, we have:


"T=2\pi\sqrt{\dfrac{L}{g}},""g=\dfrac{4\pi^2L}{T^2},""g=\dfrac{4\pi^2\cdot1.5\ m}{(3.7\ s)^2}=4.32\ \dfrac{m}{s^2}."